OpenStudy (anonymous):

Please help! how do you solve log base4 (2x+1) = log base4 (x+2) - log base4 3

OpenStudy (anonymous):

\[\log_{4} \left( 2x+1 \right)=\log_{4} \left( x+2 \right)-\log_{4} 3=\log_{4} \frac{ x+2 }{ 3 }\] \[2x+1=\frac{ x+2 }{ 3 },solve~ for~ x\]

OpenStudy (anonymous):

Do i put 2x +1 over 1 and cross multiply?

OpenStudy (anonymous):

correct.

OpenStudy (anonymous):

ok im going to do that now thank u!

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

i got negative 1/5 as my answer? is that correct @surjithayer

OpenStudy (anonymous):

correct.

OpenStudy (anonymous):

thanks again!